FEL Upgrade Project Weekly Brief
September 4 - 8, 2000

Management Highlights:

A presentation on biological applications of the FEL was given to the Virginia Joint (House/Senate) Subcommittee on Science and Technology.  The group is investigating ways the state legislature can encourage the development of biotechnology in Virginia.   The role of the FEL as an attraction to the “superstars” of the field was discussed.
A small group visited Fishkill, NY, to investigate the compatibility of the IR Upgrade design with higher energy photon production.  Physical interferences, control system status, and documentation were all reviewed.  No major hurdles were seen.
Design/Construction Activities:

Beam Physics (WBS 3.0):

WBS 3 activities continued apace this week. A JLAB Technical Note documenting longitudinal phase space management issues in the Upgrade driver is forthcoming, CSR work is ongoing, and discussions/work on BBU/other wakefield, space charge, and lattice error effects are underway.

WBS 4.0

We installed a nitrogen implanted SST cathode with a 1 um polished SST anode into the field emission test system. We then started HV conditioning with a 1 cm gap up to 105 kV before baking to ensure the system is operating properly.  Work will proceed on this next week.

 RF Systems (WBS 6):

 Writing continued on the specifications for the cathode power supply.
Assembly continued on the Zone 3 and Zone 4 Rf systems.

Instrumentation (WBS 8):

A technician to support development of the instrumentation and controls was chosen for hire.  The offer package is being assembled.
We continued preparing viewer and BPM bid packages and resolved several issues regarding procurement approach, make/buy, etc.

Beam Transport System (WBS 9)

 We started detailing the GX Prototype.
 We have found the optimum bump on the pole edge of the GX dipole using a 2D magnetic model.  We are now introducing that bump into the 3D magnetic model.
 Poles:  L&L is setting up fixturing for machining of first article quadrant.  Plan on visiting vendor today.
       Coils:  The fifth (spare) coil was confirmed DOA, lacking ~ 50 of the 350 turns.  There are no plans to build a replacement.  J. Dail has completed fabrication of the cooling plates with the attached cooling lines.  He is in the process of prototyping the process of attaching the cooling plates onto the coils.
 Budgetary Quote: The drawing package for the quad was sent to Everson Electric on August 29th. Expecting response by the week of September 18th.
 Business and technical questions about the proposals were sent to the remaining Offerors in the engineering services contract.

Wiggler (WBS 10)

 Manifold fabrication continues.
 We measured the pole tip gaps.  They have about a 0.002 inch gradient and the center gaps were larger than the ends by several thousandths.  Kevin is working on improving those readings and making them reproducible. An assembly procedure guaranteeing better gap uniformity is being developed.

Dispersion Section
  The specifications were signed and the package was given to the fabrication office to get bids from 6 vendors.
The coil specification and the dispersion section assembly specification were signed off.

Optics (WBS 11):

Conceptual designs of the Optical beam Position Monitoring (OBPM) system were reviewed and Okayed.  The system will be tested in the existing Optical Transport System.  Monitoring detectors have been received and position controllers of several types will be ordered, tested, and evaluated as to compatibility and reliability.
Pyroelectric array diagnostics for spectral measurements were installed in the OCR for characterization.  Initial use will be manual with the full software suite added when functionality has been confirmed.
A  pyroelectric detector with chopper was added to the OTS as a real time macropulse power diagnostic when married with a new sample and hold circuit.  This will be tested on the next laser run.
The Modemaster system was realigned and checked out in preparation for more detailed studied of mirror heating distortion and its effect on the FEL lasing process.
A designer is preparing drawings of the planned mirror test stand which will be used to characterize performance of the controllable curvature mirror.
Measurements of the HVAC and sprinkler systems uncovered potential conflicts with positioning of the new OTS.  We will decide soon which system to move.

Beam Operations
The electron beam was turned on in preparation for a User run next week.  A number of hardware problems were uncovered: a shorted transformer in the rf buncher supply, a 2.5 W amplifier blown, access control problems in the drive laser clean room, and a burned out transistor in the drive laser mode locker chassis.  As of this writing these systems were being put back together and prepared for turn on.

Privacy and Security Notice